# Template Free Reconstruction of Human-object Interaction with Procedural Interaction Generation

### 12 Dec 2023 v1

@@ A.2 HDM: Hierarachical Diffusion Model
A.3 Dataset statistics
We generate our ProciGen dataset based on interactions from BEHAVE [7] and InterCap [33], human scans from MGN [4], object shapes from ShapeNet [12], Objaverse [19] and ABO [15]. In total we rendered 1.1M interaction images with 21555 different object shapes. The distribution for object shapes and interactions per category are shown in Fig. 7 and Fig. 8. Our dataset has very diverse object shapes, especially for chairs and tables whose geometry also varies a lot in reality. Our procedural generation method is a scalable solution and it allows for generating large-scale interaction datasets with great amount of variations which is not obtainable via capturing real data.
Appendix B Additional Experiments
B.1 Analysis of $T_{0}$ for our HDM
B.2 Shape fidelity
@@
B.3 Interaction semantics
Our method predicts the segmentation of human and object, allowing separate manipulation which is important for downstream applications. To demonstrate this, we use Text2txt [13] to generate textures for the meshes extracted from $\text{PC}^{2}$ and our predicted point clouds. We show the reconstruction and generated textures in Fig. 11. It can be seen that $\text{PC}^{2}$ predictions are noisy and it does not reason human and object separately. This leads to low-quality mesh and generating coherent texture for a combined mesh of human and object is difficult. On the contrary, our method separate human and object while also predicting high quality individual shapes. This allows generating high quality texture and changing textures for human and object differently.
@@ B.3 Interaction semantics
B.4 More generalization results

### 06 Apr 2024 v3

@@ B.3 Analysis of $T_{0}$ for our HDM
B.4 Shape fidelity
@@
Figure 12: Out of distribution generalization. Our method can reconstruct some categories that are unseen in training data.
B.5 Interaction semantics
Our method predicts the segmentation of human and object, allowing separate manipulation which is important for downstream applications. To demonstrate this, we use Text2txt [13] to generate textures for the meshes extracted from $\text{PC}^{2}$ and our predicted point clouds. Other methods such as Paint-it [106] are also applicable here. We show the reconstruction and generated textures in Fig. 13. It can be seen that $\text{PC}^{2}$ predictions are noisy and it does not reason human and object separately. This leads to low-quality mesh and generating coherent texture for a combined mesh of human and object is difficult. On the contrary, our method separate human and object while also predicting high quality individual shapes. This allows generating high quality texture and changing textures for human and object differently.
@@ B.5 Interaction semantics
B.6 More generalization results
@@
For quantitative evaluation, we focus on 15 object categories that are seen from our synthetic data (Tab. 6). We test our method on three additional categories from BEHAVE and InterCap that are unseen and have GT data. The F-scores (human/object/combined) are: 0.465/0.453/0.479 (tennis racket), 0.333/0.332/0.361 (yoga mat), 0.375/0.305/0.360 (umbrella), 0.353/0.443/0.420 (all seen categories). We also show example reconstructions in Fig. 12. Our method can reconstruct unseen categories.